### v2

# Description of the structure of DG–algebras via characters on a groupoid

Abstract
We construct a description of graded derivations in group algebras. Using this result for arbitrary graduation of the group algebra, we describe all possible structures of DG algebras. The corresponding examples are given. The description is given in terms of characters on a groupoid analogous to the groupoid of inner action.
Introduction
In several previous papers one offers A.S. Mishchenko and others [AMS17, AM19, AAS23], was presented a method for studying the derivations of group algebras using the character space on the groupoid of conjugacy action. In the present paper, a similar groupoid and characters on it are constructed in order to describe graded derivations on graded group algebra and all possible structures of DG algebras on group algebras. [^1] The results of Sections 1.1 and 1.2 were obtained by A. Arutyunov, the other results were obtained jointly. The work of A. Arutyunov was supported by the RSF grant No. 25-11-00018.
DG algebra and graded derivations are frequently encountered. See e.g. [Orl23] where they are studied in the context of algebraic geometry. The independent study of graded derivations is an even better known question, see e.g. [Mei03]. The classification of possible DG strings has also been the focus of study [Mao11, MWZ22].
The main object of our study will be graded group algebra $\mathbb{C}[G]\cong\bigoplus A_{i}$ and its graded derivation. That is, such linear maps $d:\mathbb{C}[G]\rightarrow\mathbb{C}[G]$ for which holds the graded Leibniz rule
$d(uv)=d(u)v+(-1)^{|u|}ud(v),\quad\forall u,v\in A_{i}.$
In Section 1 we present a structure of a grouppod which is connected to the graded algevra structure. We will give a description of graded derivations as localy finite characters on this grouppoid (see Theorem 1.1) In Section 2, we describe the inner, quasi-inner, and central derivations for the graded case. Using this construction in the case of DG–algebras, in Section 3 we obtain a description of the structure of the DG–algebra in terms of characters and an isomorphism criterion between two DG–algebras. In addition, at the end of Section 3, we give several specific examples of graduations on group algebras and derivations that define the structure of the DG–algebra on them. The definitions and properties from combinatorial group theory that are necessary for the paper can be found in [LS01]
1 Groupoid, characters and graded derivations.
Let’s start with the definitions and fix the notation for future reference. We will study derivations over group algebras with a DG structure. A group algebra is defined in standard form as follows:
Definition 1.1.
$K[G]$ – is the group algebra of a group $G$ over a field $K$ is an associative algebra over a field $K$, the elements of which are all possible formal finite sums of the form $\sum_{g\in G}k_{g}g,$ $g\in G,$ $k_{g}\in K$, and the operations are determined by the formulas:
$\sum_{g\in G}a_{g}g+\sum_{g\in G}b_{g}g=\sum_{g\in G}(a_{g}+b_{g})g,$
$(\sum_{g\in G}a_{g}g)(\sum_{g\in G}b_{g}g)=\sum_{h\in G}(\sum_{xy=h\in G}(a_{x}b_{y})h).$
The elements of the group $G$ form the basis of the algebra $K[G]$. The multiplication of the basic elements in the group algebra is induced by the multiplication in group.
Furthermore, we need the concept of a groupoid. Typically, this is a category in which every morphisms is invertible. More precisely, we will work with the following construction associated with a finitely generated group $G$, so this category is small (objects and morphisms are sets).
Definition 1.2.
A groupoid $\Gamma$ contains a set of objects $Obj(\Gamma)$ and morphisms $Hom(\Gamma)$. Each morphism has a source $s(\phi)$ and a target $t(\phi)$, i.e. each morphism is in the set $Hom(s(\phi),t(\phi))$. Composable morphisms have an associative operation (composition). The following conditions hold:
  1. 1.
    For each object $a\in Obj(\Gamma)$, there is a neutral endomorphism $1_{a}\in Hom(a,a)$ such that $\forall\phi\in Hom(b,a)$ $1_{a}\circ\phi=\phi$ and $\forall\phi\in Hom(a,b)$ $\phi\circ 1_{a}=\phi$.
  2. 2.
    There is an inverse for each morphism. I.e. $\forall\phi\in Hom(a,b)$ exists $\psi\in Hom(b,a)$ such that $\phi\circ\psi=1_{b}$ and $\psi\circ\phi=1_{a}$.
1.1 Groupoid of conjugacy action.
The following construction is a generalization of the structure from [AMS17]. Suppose we have a graded group algebra $\mathbb{C}[G]\cong\bigoplus A_{i}$, here $i\in\mathbb{Z}$. Group $G$ assumes to be finitely generated. We will define the groupoid of conjugacy action, denoted by $\Gamma$, for the group $G$ in the sense of Definition 1.2.
[…]
Consider two morphisms $\phi=(u_{1},v_{1})$ and $\psi=(u_{2},v_{2})$ such that $t(\phi)=s(\psi)$, that is, the composition is defined for them $\psi\circ\phi$. Define the composition of morphisms as follows
$(u_{2},v_{2})\circ(u_{1},v_{1}):=((-1)^{|v_{2}|}v_{2}u_{1},v_{2}v_{1})).$
This formula is conveniently recorded in the form of a diagram.
$(-1)^{|v_{1}|}v_{1}^{-1}u_{1}=u_{2}v_{2}^{-1}$$u_{1}v_{1}^{-1}$$(-1)^{|v_{2}|}v_{2}^{-1}u_{2}$$\scriptstyle{(u_{2},v_{2})}$$\scriptstyle{(u_{1},v_{1})}$$\scriptstyle{(u_{2},v_{2})\circ(u_{1},v_{1})}$
Proposition 1.1.
$\Gamma$ is a groupoid.
Proof.
First of all, we note that $\exists g\rightarrow h\Leftrightarrow\exists t:h=(-1)^{|t|}tgt^{-1}.$
Let we have $\phi=(u_{1},v_{1})$ and $\psi=(u_{2},v_{2})$ such that $t(\phi)=s(\psi)$. Let’s say $s(\phi)=a$, $t(\phi)=s(\psi)=b$ and $t(\psi)=c$.
Check that
$s(\psi\circ\phi)=a$ and $t(\psi\circ\phi)=c$.
By definition
$s(\psi\circ\phi)=(-1)^{|v_{2}v_{1}|}v_{1}^{-1}v_{2}^{-1}(-1)^{|v_{2}|}v_{2}u_{1}$ = $(-1)^{|v_{1}|}v_{1}^{-1}u_{1}=a$.
Calculate that
$t(\psi\circ\phi)$ = $(-1)^{|v_{2}|}v_{2}u_{1}v_{1}^{-1}v_{2}^{-1}=u_{2}v_{2}^{-1}=c$, because $u_{1}v_{1}^{-1}=(-1)^{|v_{2}|}v_{2}^{-1}u_{2}$
Now we prove the associativity of the composition. Let $\phi_{1},\phi_{2},\phi_{3}$ - morphisms such that $s(\phi_{1})=t(\phi_{2})$ and $s(\phi_{2})=t(\phi_{3})$, then
$\phi_{1}\circ(\phi_{2}\circ\phi_{3})=(\phi_{1}\circ\phi_{2})\circ\phi_{3}.$
We proved above that all compositions are defined correctly. Let’s calculate the left part.
$\phi_{2}\circ\phi_{3}=((-1)^{|v_{2}|}v_{2}u_{3},v_{2}v_{3}))$ $\phi_{1}\circ(\phi_{2}\circ\phi_{3})=(u_{1},v_{1})\circ((-1)^{|v_{2}|}v_{2}u_{3},v_{2}v_{3}))=((-1)^{|v_{1}|+|v_{2}|}v_{1}v_{2}u_{3},v_{1}v_{2}v_{3}).$
Now let’s calculate the right side of the expression
$(\phi_{1}\circ\phi_{2})\circ\phi_{3}=((-1)^{|v_{1}|}v_{1}u_{2},v_{1}v_{2})\circ(u_{3},v_{3})=((-1)^{|v_{1}|+|v_{2}|}v_{1}v_{2}u_{3},v_{1}v_{2}v_{3}).$
The left and right sides of the expression are equal, which means that the operation $\circ$ is associative.
Let’s show that there is an element $1_{g}$, such that:
$s(1_{g})=t(1_{g})=g$
$\phi\cdot 1_{s(\phi)}=\phi=1_{t(\phi)}\cdot\phi$, $\forall\phi\in$ $Hom(\Gamma)$
Let’s take as a $1_{g}$ element $(g,e)$, where e is a trivial element of the group G.
$(-1)^{|e|}e^{-1}g=g=ge^{-1}$$\scriptstyle{(g,e)}$
Indeed, let’s take an arbitrary morphism $(u,v)$, than:
$(uv^{-1},e)(u,v)=(u,v)$.
$(u,v)((-1)^{|v|}v^{-1}u,e)=((-1)^{|v|}v(-1)^{|v|}v^{-1}u,e)=(u,v)$.
We have shown that $\Gamma$ is a category. Now it is necessary to check the inversibility of the arrows.
We show the presence of an inverse morphism $\forall\phi\in Hom(\Gamma)$. For morphism of the form (u,v) let’s call the inverse morphism $((-1)^{|v|}v^{-1}uv^{-1},v^{-1})$.
Indeed:
$(u,v)((-1)^{|v|}v^{-1}uv^{-1},v^{-1})=(uv^{-1},e)=(t(\phi),e).$
$((-1)^{|v|}v^{-1}uv^{-1},v^{-1})(u,v)=((-1)^{|v|}v^{-1}u,e)=(s(\phi),e).$
$(-1)^{|v|}v^{-1}u$$uv^{-1}$$\scriptstyle{(u,v)}$$\scriptstyle{((-1)^{|v|}v^{-1}uv^{-1},v^{-1})}$
Thus, the Proposition 1.1. is proved. ∎
The following properties of the groupoid are easily verified by direct calculation.
Proposition 1.2.
For a groupoid $\Gamma$ holds:
  1. 1.
    For endomorphisms $Hom(a,a),a\in\pm G$ holds the following descriptions
    $Hom(a,a)=\{\chi\in Hom(\Gamma)\mid s(\chi)=(-1)^{|t|}t^{-1}a,$ $t(\chi)=at^{-1},$ $t\in Z(a)$ and $t\in$ $A_{2i}[G]\}$
  2. 2.
    For $a]:=\{b\in\pm G\mid\exists t\in\pm G:b=(-1)^{|t|}t^{-1}at\}$ denote by $\Gamma_{[a]}$, a subgroupoid such that $Obj(\Gamma_{[a]})=[a]$. The following properties holds:
    1. (a)
      $Hom(\Gamma_{[a]})=\{\phi\in Hom(\Gamma)\mid s(\phi)\in[a],t(\phi)\in[a]\}$.
    2. (b)
      $\Gamma=\coprod\Gamma_{[a]}$.
  3. 3.
    There is a left action $Hom(a,a)\times Hom(a,b)\rightarrow Hom(a,b)$, $(\phi,\psi)\to\psi\circ\phi$. This action is free and transitive.
Let us show that with the help of characters of $\Gamma$, it is possible to describe a derivation satisfying the graded Leibniz rule.
Definition 1.3.
We say that the linear operator $d$: $\mathbb{C}[G]$ $\to$ $\mathbb{C}[G]$ satisfies the graded Leibniz rule if
$d(uv)=d(u)v+(-1)^{|u|}ud(v)$, $\forall u,v\in A_{i}$
Definition 1.4.
We call the map $\chi$ : $Hom(\Gamma)$ $\to$ $\mathbb{C}$ a character if it satisfies the identity $\chi(\psi\circ\phi)=\chi(\psi)+\chi(\phi)$
$\bullet$$\bullet$$\bullet$$\scriptstyle{\chi(\psi)}$$\scriptstyle{\chi(\phi)}$$\scriptstyle{\chi(\psi\circ\phi)=\chi(\psi)+\chi(\phi)}$
Definition 1.5.
We will say that the character $\chi$ is locally finite if for all $v\in pmG$ we get that $\chi(u,v)=0$ for all except a finite number $u$.
We will denote by $\chi(\Gamma)$ the space of local finite characters. In section 1.2 we will show that $\chi(\Gamma)$ has a natural structure of Lie algebra.
The graded Leibniz rule in a group algebra can be written in terms of characters on the groupoid $\Gamma$ as follows.
Proposition 1.3.
For any graded derivation $d$ there exists a locally finite character such that
$d(u)=\sum\limits_{g\in G}\Big{(}\sum\limits_{h\in\pm G}\chi(\phi(h,g)\lambda^{h}\Big{)}g,\forall u\in A_{i}$ (1)
Proof.
It is enough to check the statement on the elements $g\in\pm G$, since they generate $\mathbb{C}[G]$, as a vector space.
For any element $g$ of group G we have $d(g)=\sum\limits_{h\in G}d_{g}^{h}h$
Define $\chi(\phi(h,g)):=d_{g}^{h}$, $\chi(\phi(-h,g)):=-d_{g}^{h}$ and let us show that it is a character. We have
$d(g_{2}g_{1})=d(g_{2})g_{1}+(-1)^{|g_{2}|}g_{2}d(g_{1}).$
Let’s rewrite this in a new notation:
$\sum\limits_{h\in G}d_{g_{2}g_{1}}^{h}h=\sum\limits_{h\in G}d_{g_{2}}^{h}hg_{1}+(-1)^{|g_{2}|}g_{2}\sum\limits_{h\in G}d_{g_{1}}^{h}h.$
It is clear that summation over all $h$ from the group for $h$ and $gh$ is equivalent, so we can convert the right-hand sides to the following form:
$\sum\limits_{h\in G}d_{g_{2}g_{1}}^{h}h=\sum\limits_{h\in G}d_{g_{2}}^{hg_{1}^{-1}}h+\sum\limits_{h\in G}d_{g_{1}}^{(-1)^{|g_{2}|}g_{2}^{-1}h}h\Rightarrow d_{g_{2}g_{1}}^{h}=d_{g_{2}}^{hg_{1}^{-1}}+d_{g_{1}}^{(-1)^{|g_{2}|}g_{2}^{-1}h}.$ (2)
Let’s make a replacement:
$h_{1}=hg_{1}^{-1},$
$h_{2}=(-1)^{|g_{2}|}g_{2}^{-1}h.$
Then the last equality (2) takes the form:
$d_{g_{2}g_{1}}^{(-1)^{|g_{2}|}g_{2}h_{2}}=d_{g_{2}}^{h_{1}}+d_{g_{1}}^{h_{2}}$ $\Leftrightarrow$ $\chi(\phi((-1)^{|g_{2}|}g_{2}h_{2},g_{2}g_{1}))=\chi(\phi(h_{1},g_{2}))+\chi(\phi(h_{2},g_{1})).$
It remains only to note that $((-1)^{|g_{2}|}g_{2}h_{2},g_{2}g_{1})=(h_{1},g_{2})(h_{2},g_{1}).$
A corollary of the above statement is the following theorem, the proof of which is similar to the non-graded case, see Theorem 1 of [Aru20].
Theorem 1.1.
For any graded derivation $d$ there exists such a unique locally finite character of $\chi_{d}$, that the formula is valid
$d(a)=(-1)^{|a|}a\Big{(}\sum\limits_{t\in G}\chi((-1)^{|a|}at,a)t\Big{)}.$
In fact, we can say that we have an isomorphism $Der(\mathbb{C}[G])\cong\chi(\Gamma)$.
To prove this, we need to use formula (1) to element $a\in\pm G$ and write an arbitrary element $g\in G$ as $g=at$ and then remove the terms with a zero coefficient from the sum. See the proof of Theorem 1 in [Aru20] for more details.
1.2 Characters.
Locally finite characters are isomorphic to derivations, which means that the binary commutator operation on derivations generates a binary operation on characters. Formula (1) allows us to define a composition in character space. Let $d_{1},d_{2}$ be the derivations, defined by the characters $\chi_{d_{1}},\chi_{d_{2}}$. The character that corresponds to the commutator of derivations $d_{1},d_{2}]$ is denoted by $\chi_{[d_{1},d_{2}]}$. It defines an operation on the character space.
$\{\chi_{d_{1}},\chi_{d_{2}}\}:=\chi_{[d_{1},d_{2}]}$ (3)
For this operation, we have the following equality
Proposition 1.1.
Holds the formula
$\{\chi_{d_{1}},\chi_{d_{2}}\}(a,g)=\sum\limits_{h\in G}\chi_{d_{1}}(a,h)\chi_{d_{2}}(h,g)-\chi_{d_{2}}(a,h)\chi_{d_{1}}(h,g).$ (4)
Proof.
Let $g\in G$, than we have
$d_{1}(g)=\sum\limits_{h\in G}\chi_{d_{1}}(h,g)h$,
$d_{2}(g)=\sum\limits_{h\in G}\chi_{d_{2}}(h,g)h$,
$\{\chi_{d_{1}},\chi_{d_{2}}\}(g)=\sum\limits_{h\in G}\{\chi_{d_{1}},\chi_{d_{2}}\}(a,g)a$.
Write down the equation for the commutator
$d_{1},d_{2}]=d_{1}d_{2}-d_{2}d_{1}$,
$d_{1}d_{2}(g)=\sum\limits_{h\in G}\chi_{d_{2}}(h,g)\Big{(}\sum\limits_{a\in\pm G}\chi_{d_{1}}(a,h)a\Big{)}$,
$d_{2}d_{1}(g)=\sum\limits_{h\in G}\chi_{d_{1}}(h,g)\Big{(}\sum\limits_{a\in\pm G}\chi_{d_{2}}(a,h)a\Big{)}$.
Due to local finiteness we can change the order of summation in the last equation and get
$d_{1},d_{2}](h)=\sum\limits_{a\in G}\Big{(}\sum\limits_{h\in G}\chi_{d_{2}}(h,g)\chi_{d_{1}}(a,h)-\chi_{d_{1}}(h,g)\chi_{d_{2}}(a,h)\Big{)}a$.
Equation $\{\chi_{d_{1}},\chi_{d_{2}}\}(a,g)$ it is coefficient under $a$, so
$\{\chi_{d_{1}},\chi_{d_{2}}\}(a,g)=\sum\limits_{h\in G}\chi_{d_{1}}(a,h)\chi_{d_{2}}(h,g)-\chi_{d_{2}}(a,h)\chi_{d_{1}}(h,g)$.
It follows that the pair $(\chi_{\Gamma},\{\cdot;\cdot\})$ is a Lie algebra which is isomorphic by proposition 1.3 to the algebra of all graded derivations.
2 Examples of graded derivations and their description using characters on a groupoid.
We will demonstrate the correspondence between characters on groupoid and graded derivations using examples of inner derivations, central derivations and quasi-inner derivations. It gives give easy examples of such operators.
2.1 Inner derivations
Consider the element $a\in\pm G$. Let’s define the mapping $\chi^{a}:Hom(\Gamma)\rightarrow\mathbb{C}$ as follows (assuming that $a\neq b$):
$\chi^{a}(\phi):=\begin{cases}1,&\text{$\phi\in Hom(b,a)$}\\ -1,&\text{$\phi\in Hom(a,b)$}\\ 0,&\text{else.}\end{cases}$ (5)
It is clear that $\chi^{a}$ is trivial on endomorphisms.
Proposition 2.1.
Map $\chi^{a}$ sets the inner derivations for $a\in\pm G$ by formula
$d_{a}:x\rightarrow[a,x]_{grad}$ , where
$a,x]_{grad}=ax-(-1)^{|x|}xa$.
Proof.
Let’s check that $d_{a}$ is a graded derivation (that is, satisfies the graded Leibniz rule):
$d_{a}(x)y+(-1)^{|x|}xd_{a}(y)=(ax-(-1)^{|x|}xa)y+(-1)^{|x|}x(ay-(-1)^{|y|}ya)=axy-(-1)^{|xy|}xya=d_{a}(xy)$.
Let’s check that it really sets the necessary derivation. Due to linearity, it is enough to check the statement on the element $g\in\pm G$. According to the previous corollary, we have an equality:
$d_{a}(g)=\sum\limits_{h\in G}\chi(h,g)h.$
We have that $\chi(h,g)\neq 0\Leftrightarrow t((h,g))=a$ or $s((h,g))=a$. Let’s consider these two cases:
  1. 1.
    $s((h,g))=a\Leftrightarrow(-1)^{|g|}g^{-1}h=a\Leftrightarrow h=(-1)^{|g|}ga.$
  2. 2.
    $t((h,g))=a\Leftrightarrow hg^{-1}=a\Leftrightarrow h=ag.$
Therefore, we have: $d_{a}(g)=ag-(-1)^{|g|}ga.$
2.2 Central derivations
We will now give an example of a derivation that is not inner. It will be demonstrated that such a derivation has an easy calculation form and is not trivial on endomorphisms. It will thus be established that they cannot be presented as a (perhaps formal) sum of inner derivations.
Definition 2.1.
We say that map $\tau:G\rightarrow\mathbb{C}$ – is graded group character, if it satisfies the following property:
$\tau(ab)=\tau(a)+(-1)^{|a|}\tau(b),$ $\forall a,b\in G$.
Let’s fix the graded group character $\tau$, and let $z\in Z(G)$ be an element of group center.
Definition 2.2.
Define map $d_{z}^{\tau}:\mathbb{C}[G]\rightarrow\mathbb{C}[G]$ on the generators $g\in\pm G$ as
$d_{z}^{\tau}:g\rightarrow\tau(g)gz.$
Continued by linearity for the whole algebra $\mathbb{C}[G]$ map $d_{z}^{\tau}$ we will call it as the central graded derivation.
Proposition 2.1.
The map $d_{z}^{\tau}$ is a graded derivation.
Proof.
[…]
$d_{z}^{\tau}(uv)=\tau(uv)uvz=\tau(u)uzv+(-1)^{|u|}u\tau(v)vz=d_{z}^{\tau}(u)v+(-1)^{|u|}ud_{z}^{\tau}(v)$.
On arbitrary elements of a group algebra, the graded Leibniz rule is fulfilled due to the linearity of the $d_{z}^{\tau}$. ∎
Proposition 2.2.
Nontrivial central derivations are not inner.
Proof.
Let $\chi_{\tau,z}$ – a character that corresponds to derivation $d_{z}^{\tau}$. By corollary 1 we have equality
$d_{z}^{\tau}(g)=\tau(g)gz=g\Big{(}\sum\limits_{t\in G}\chi((-1)^{|g|}gt,g)(-1)^{|g|}t\Big{)}$.
Therefore, the character is not equal to 0 only in the case of $t=z$:
$\chi((-1)^{|g|}gz,g)=(-1)^{|g|}\tau(g)$.
Denote $\phi=((-1)^{|g|}gz,g)$ and calculate its beginning and end
$s(\phi)=z$,
$t(\phi)=(-1)^{|g|}z.$
Then in the case when $g\in A_{2n}$, we have: $s(\phi)=t(\phi)$ $\Rightarrow\chi_{\tau,z}$ non trivial on a loop $\phi$, and $d_{z}^{\tau}$ it is not an inner derivation. ∎
A direct check shows that inner and central derivations are subalgebras. For more details see [Aru20].
2.3 Quasi-inner derivations
As seen from formula (5) inner derivations are trivial on endomorphisms; however, the opposite is not true. We will call the class of such derivations quasi-inner.
Definition 2.3.
Define QInnDer($\mathbb{C}[G]$) derivation as follows:
$QInnDer(\mathbb{C}[G])=\left\{d\in Der\mid\forall a\in Obj(\Gamma),\forall\phi\in Hom(a,a)\Rightarrow\chi_{d}(\phi)=0\right\}.$ (6)
Quasi-inner derivations are clearly a sub-space, but the more significant statement is that they constitute an ideal in algebra of all derivations.
Proposition 2.3.
QInnDer($\mathbb{C}[G]$) $\triangleleft$ Der($\mathbb{C}[G]$) is an ideal.
Proof.
Let’s proof that QInnDer($\mathbb{C}[G]$) is a subalgebra, i.e.
$d_{1}$, $d_{2}$ $\in$ QInnDer($\mathbb{C}[G]$) $\Rightarrow$ $\left[d_{1},d_{2}\right]\in QInnDer(\mathbb{C}[G])$.
character $\chi_{d_{1}}$ can be represented as:
$\chi_{d_{1}}=\sum\limits_{a\in G}\lambda^{a}\chi^{a}$, $\lambda^{a}\in\mathbb{C}$,
since $\chi_{d_{1}}$ is trivial on endomorphisms, where $\chi_{d_{1}}$ defined by formula (5). Similarly with a character $\chi_{d_{2}}$:
$\chi_{d_{2}}=\sum\limits_{b\in G}\lambda^{b}\chi^{b}$, $\lambda^{b}\in\mathbb{C}$.
Because of the bilinearity of the commutator,
$\{\chi_{d_{1}},\chi_{d_{2}}\}=\sum\limits_{a\in G}\sum\limits_{b\in G}\lambda^{a}\mu^{b}\{\chi^{a},\chi^{b}\}$.
The commutator of the $d^{a}$ и $d^{b}$ can be represented as follows:
$d^{a},d^{b}]=d^{ab}-d^{ba}$.
Then define the character $\{\chi^{a},\chi^{b}\}$ by the following formula
$\{\chi^{a},\chi^{b}\}=\chi^{ab}-\chi^{ba}$.
Obtain the final formula for $\{\chi_{d_{1}},\chi_{d_{2}}\}$:
$\{\chi_{d_{1}},\chi_{d_{2}}\}=\sum\limits_{a\in G}\sum\limits_{b\in G}\lambda^{a}\mu^{b}\chi^{ab}-\sum\limits_{a\in G}\sum\limits_{b\in G}\lambda^{a}\mu^{b}\chi^{ba}$,
$\{\chi_{d_{1}},\chi_{d_{2}}\}\in QInnDer(\mathbb{C}[G])$, to proof that, consider the value of the character on the loop $(uz,z),z\in Z_{G}(u)$
$\{\chi_{d_{1}},\chi_{d_{2}}\}(uz,z)=\sum\limits_{ab=zuz^{-1}}\lambda^{a}\mu^{b}-\sum\limits_{ab=u}\lambda^{a}\mu^{b}-\sum\limits_{ba=u}\lambda^{a}\mu^{b}+\sum\limits_{ba=zuz^{-1}}\lambda^{a}\mu^{b}=0$.
Now pass directly to the proof of the theorem. Represent $\chi_{d_{0}}$ as
$\chi_{d_{0}}=\sum\limits_{a\in G}\lambda^{a}\chi^{a}$.
Prove the theorem for $\chi^{a}$ and extend the result to the $\chi_{d_{0}}$, using the bilinearity of the commutator. Consider character $\{\chi_{d},\chi^{a}\}$. Proof that it is trivial on endomorphisms, i.e. $\forall b\in G$ and for $\forall z\in Z_{G}(b)$ is performed as $\{\chi_{d},\chi^{a}\}(bz,z)=0$. By Proposition 3.1:
$\{\chi_{d},\chi^{a}\}(bz,z)=\sum\limits_{h\in G}\chi_{d}(bz,h)\chi^{a}(h,z)-\chi^{a}(bz,h)\chi_{d}(h,z)$.
$\chi^{a}(h,z)\neq 0$ only in 2 cases: when $h=za$ and $h=az$. $\chi^{a}(bz,h)\neq 0$ only in 2 cases: when $h=bza^{-1}$ and $h=a^{-1}bz$. It means that
$\{\chi_{d},\chi^{a}\}(bz,z)=\chi_{d}(bz,za)-\chi_{d}(bz,az)+\chi_{d}(a^{-1}bz,z)-\chi_{d}(bza^{-1},z)$.
But also
$(a^{-1}bz,z)\circ(bz,za)=(bz,az)\circ(bza^{-1},z)$.
That means
$\chi_{d}(bz,za)+\chi_{d}(a^{-1}bz,z)=\chi_{d}(bz,az)+\chi_{d}(bza^{-1},z)$,
$\{\chi_{d},\chi^{a}\}(bz,z)=0$.
It is easy to see that central derivations are not quasi-inner, for which it is sufficient to repeat the reasoning from proposition 2.2.
3 Description of the structure of a DG–algebra in terms of characters on a groupoid.
Our next goal is to give the description of possible DG–algebra sctructure up to isomorphism. We follow the definitions from [FHT01]
Definition 3.1.
Differential graded algebra (DG–algebra) is a graded algebra $A\cong\bigoplus A_{i}$, equipped with derivation
$d:A\rightarrow A$,
which has a degree of $1$ (the case of a cochain complex) or 1-1 (the case of a chain complex) and satisfies following properties
  1. 1.
    $d\circ d=0$,
  2. 2.
    $d(uv)=d(u)v+(-1)^{|u|}ud(v)$.
We will give the necessary and sufficient conditions from the point of view of characters on a groupoid that derivatiation sets the structure of a DG–algebra on a graded group algebra. We will describe the case of a cochain complex, since the reasoning is similar for a chain complex.
Theorem 3.1.
The characters on the groupoid define the structures of the DG–algebra if and only if the following conditions are satisfied:
  1. 1.
    The convolution is trivial:
    $\sum\limits_{h\in G}\chi(h,g)\chi(x,h)=0,\quad\forall g,x\in G.$ (7)
  2. 2.
    For all $(h,g)\in Hom(\Gamma_{A_{i}})$, $i\neq 1$, we have $\chi(h,g)=0$.
The main idea is to prove that the first condition is equivalent to $d^{2}(g)=0$, and that the second condition is equivalent to $d(A^{i})\subseteq A^{i+1}$.
Proof.
Using proposition 2.2, we rewrite the properties of the DG–algebra in terms of characters.
  1. 1.
    By direct calculation we can see that
    $d^{2}(g)=d\Big{(}\sum\limits_{h\in G}d^{h}_{g}h\Big{)}=\sum\limits_{h\in G}d^{h}_{g}\Big{(}\sum\limits_{x\in G}d^{x}_{h}x\Big{)}=\sum\limits_{h\in G}(h,g)\Big{(}\sum\limits_{x\in G}(x,h)x\Big{)}=\sum\limits_{x\in G}\Big{(}\sum\limits_{h\in G}(\chi(h,g)\chi(x,h))\Big{)}x.$
    So the condition (7) holds.
  2. 2.
    Let $g\in A^{i}$, than $d(g)=\sum\limits_{h\in G}(h,g)h\in A^{i+1}\Leftrightarrow(h,g)=0$, if $h\notin A^{i+1}$.
    Note that if $u\in A^{i+1},v\in A^{i}$, than for $\phi=(u,v)$ we have: $s(\phi)=(-1)^{|v|}v^{-1}u\in A^{1},t(\phi)=uv^{-1}\in A^{1}$,this means that in the case of a DG algebra, the characters may not be equal to 0 only on morphisms between elements from $A^{1}$.
Now let us recall the definition of an isomorphism with respect to the structure of DG–algebra.
Definition 3.2.
An isomorphism $f$ between two DG–algebras $(A_{1},d_{1})$ and $(A_{2},d_{2})$, $f$: $A_{1}$ $\to$ $A_{2}$, is an algebra isomorphism such that following properties hold:
  1. 1.
    f holds the structure of graded algebra $f(A^{i}_{1})\subset A^{i}_{2}$.
  2. 2.
    for all $a\in A_{1}$ holds $f(d_{1}(a))=d_{2}(f(a))$.
Now we will describe the structures of DG–algebras up to isomorphism from definition (3.2).
Let $X_{i},i=1,2$ be a matrix (if $G$ is an infinite group, that this matrix is also infinite) in following sense
$X:G\times G\to\mathbb{C},$
$(x,h)\mapsto\chi(x,h).$
Theorem 3.2.
Two DG–algebras $(\mathbb{C}[G],d_{1})$ and $(\mathbb{C}[G],d_{2})$ are isomorphic iff there is a matrix C, such that
  1. 1.
    matrixes $X_{1}$ and $X_{2}$ are conjugated
    $CX_{2}=X_{1}C,$
  2. 2.
    The mapping f is $Aut(\mathbb{C}[G])$.
In other words that means that sets of constants $c_{h}^{g}$, gives $f(g)=\sum\limits_{h\in G}c_{h}^{g}h\in Aut(\mathbb{C}[G])$ and
$\sum\limits_{h\in G}c_{h}^{g}\chi_{2}(x,h)=\sum\limits_{h\in G}\chi_{1}(h,g)c_{x}^{h},\forall g,x\in G,$ (8)
where $\chi_{1}$ and $\chi_{2}$ denotes characters that define derivations $d_{1}$ and $d_{2}$.
Proof.
Suppose that we have $f:(\mathbb{C}[G],d_{1})\rightarrow(\mathbb{C}[G],d_{2})$ — an isomorphism of DG–algebras given on generators by the formula $f(g)=\sum\limits_{h\in G}c_{h}^{g}h$. By assumption we have:
$f(d_{1}(g))=d_{2}(f(g))$.
Then
$f(d_{1}(g))=f(\sum\limits_{h\in G}\chi_{1}(h,g)h)=\sum\limits_{h\in G}\chi_{1}(h,g)f(h)=\sum\limits_{h\in G}\chi_{1}(h,g)(\sum\limits_{x\in G}c_{x}^{h}x)=(\sum\limits_{h\in G}\chi_{1}(h,g)c_{x}^{h})x$;
$d_{2}(f(g))=d_{2}(\sum\limits_{h\in G}c_{h}^{g}h)=\sum\limits_{h\in G}c_{h}^{g}d_{2}(h)=\sum\limits_{h\in G}c_{h}^{g}(\sum\limits_{x\in G}\chi_{2}(x,h)x)=(\sum\limits_{h\in G}c_{h}^{g}\chi_{2}(x,h))x$.
From where we get our equality (8). To prove the converse, just define $f(g)=\sum\limits_{h\in G}c_{h}^{g}h$ and for the same reasons, we get the necessary equalities. ∎
Corollary 3.1.
In the case when $f\in Aut(G)$ equality (8) takes the form
$\chi_{1}(h,g)=\chi_{2}(f(h),f(g)).$
That is, all characters giving isomorphic structures on a group algebra differ by the action of the automorphism $f$ on the groupoid $\Gamma$.
Example 3.1.
In group algebra $\mathbb{C}[G]\cong\bigoplus A^{i}$, such that $A^{1}\neq 0$ and $\exists a\in Z(G)$ and $a\in A^{1}$, the inner derivation generated by the element $a$ according to the formula from proposition 4.1 defines the structure of the DG–algebra on $\mathbb{C}[G]$. Note that identity element of group lies in $A^{0}$.
Proof.
It must be shown that $d^{2}=0$ is fulfilled for such derivations. We know from Theorem (3.1) that this condition is equivalent to
$\sum\limits_{h\in G}\chi(h,g)\chi(x,h)=0$, $\forall g,x\in G.$
Consider the product $\chi(h,g)\chi(x,h)$, by definition of inner derivations, this product is not equal to 0 if and only if $(h,g)$ and $(x,h)$ $\in Hom(a,-a)$ or Hom(a,a)Hom(-a,a).
Let $s(h,g)=s(x,h)=a$ and t(h,g)=t(x,h)=at(h,g)=t(x,h)=-a, then $(-1)^{|g|}g^{-1}h=(-1)^{|h|}h^{-1}x$ and $hg^{-1}=xh^{-1}$, but it is impossible because $a\in A^{1}\Rightarrow|h|\neq|g|.$
Let $s(h,g)=t(x,h)=a$ and t(h,g)=s(x,h)=at(h,g)=s(x,h)=-a, then $(-1)^{|g|}g^{-1}h=xh^{-1}$ and $hg^{-1}=(-1)^{|h|}h^{-1}x,$ but this is similarly impossible because $|h|\neq|g|.$
The remaining two cases are similar. ∎
Proposition 3.1.
Let $d^{\tau}_{z}(g)=\tau(g)gz$ - the central derivation given by the element $z\in A^{1}$. It defines the structure of a DG–algebra if and only if $\tau(g)\tau(gz)=0,\forall g\in G$.
Proof.
The fact that $d^{\tau}_{z}(g):A^{i}\rightarrow A^{i+1}$ follows from the fact that $z\in A^{1}$.
It remains to check under what conditions $d^{\tau}_{z}(g)^{2}=0$. To do this, we simply write by definition:
$(d^{\tau}_{z})^{2}(g)=d^{\tau}_{z}(\tau(g)gz)=\tau(g)\tau(gz)gz^{2}\Rightarrow(d^{\tau}_{z})^{2}(g)=0\Leftrightarrow\tau(g)\tau(gz)=0.$
The example below illustrates a non-trivial condition that is implied by the previous proposition.
Example 3.2.
[…]
Sets the structure of the DG–algebra on $\mathbb{C}[G]$.
[…]
Proof.
To prove, it is necessary to check only the fulfillment of the graduated Leibniz rule. And to do this, we need to make sure that the given $\tau$ is really a group character and take advantage of Proposition (3.1). It is not difficult to do this by simply going through all possible options. ∎
Corollary 3.1.
In group graded algebra $\mathbb{C}[Z]\cong\bigoplus\mathbb{C}[x^{k}]$, where $\mathbb{Z}=\left\langle{x}\right\rangle$, it is possible to define the structure of the DG algebra using central derivation $d^{\tau}_{x}$, where
[…]
Corollary 3.2.
Consider the Heisenberg integer group of 3×3 upper triangular matrices
$H_{3}=\left\langle{x,y,z\mid z=xyx^{-1}y^{-1},xz=zx,yz=zy}\right\rangle$.
It is well known that any element of this group can be represented as $y^{b}z^{c}x^{a}$. Then $f(y^{b}z^{c}x^{a})=c\in\mathbb{Z}$ - an epimorphism into a group of integers. From here we get the graduation $\mathbb{C}[H_{3}]\cong\bigoplus_{k\in\mathbb{Z}}\mathbb{C}[z^{k}\left\langle{x,y}\right\rangle]$. And the central differentiation $d^{\tau}_{z}$, constructed similarly to Example 5.2, gives the structure of the DG–algebra on $\mathbb{C}[H_{3}]$.

### v3

@@
-Description of the structure of DG–algebras via characters on a groupoid
+Description of DG–algebras structure via characters on a groupoid
 Abstract
 We construct a description of graded derivations in group algebras.
-Using this result for arbitrary graduation of the group algebra, we describe all possible structures of DG algebras.
-The corresponding examples are given.
-The description is given in terms of characters on a groupoid analogous to the groupoid of inner action.
+Using this result for arbitrary grading of the group algebra, we describe all possible DG–algebras structures.
+Examples are given.
+The description is given in terms of characters on a groupoid analogous to the groupoid of conjugate action.
 Introduction
-In several previous papers one offers A.S.
+In several previous papers bye A.S.
 Mishchenko and others [AMS17, AM19, AAS23], was presented a method for studying the derivations of group algebras using the character space on the groupoid of conjugacy action.
-In the present paper, a similar groupoid and characters on it are constructed in order to describe graded derivations on graded group algebra and all possible structures of DG algebras on group algebras.
+In the present paper, a similar groupoid and characters on it are constructed in order to describe graded derivations on graded group algebras and all possible DG–algebra structures on group algebras.
 [^1] The results of Sections 1.1 and 1.2 were obtained by A.
@@
 25-11-00018.
-DG algebra and graded derivations are frequently encountered.
+DG–algebras and graded derivations are occur frequently.
 See e.g.
-[Orl23] where they are studied in the context of algebraic geometry.
-The independent study of graded derivations is an even better known question, see e.g.
+[Orl23], where they are studied in the context of algebraic geometry.
+Graded derivations in isolation are an even better known subject, see e.g.
 [Mei03].
-The classification of possible DG strings has also been the focus of study [Mao11, MWZ22].
-The main object of our study will be graded group algebra $\mathbb{C}[G]\cong\bigoplus A_{i}$ and its graded derivation.
-That is, such linear maps $d:\mathbb{C}[G]\rightarrow\mathbb{C}[G]$ for which holds the graded Leibniz rule
+The classification of possible DG structures has also been a focus of study [Mao11, MWZ22].
+The main object of our study will be the graded group algebra $\mathbb{C}[G]\cong\bigoplus A_{i}$ and its graded derivations, that is linear maps $d:\mathbb{C}[G]\rightarrow\mathbb{C}[G]$ satisfying the graded Leibniz rule
 $d(uv)=d(u)v+(-1)^{|u|}ud(v),\quad\forall u,v\in A_{i}.$
-In Section 1 we present a structure of a grouppod which is connected to the graded algevra structure.
-We will give a description of graded derivations as localy finite characters on this grouppoid (see Theorem 1.1) In Section 2, we describe the inner, quasi-inner, and central derivations for the graded case.
-Using this construction in the case of DG–algebras, in Section 3 we obtain a description of the structure of the DG–algebra in terms of characters and an isomorphism criterion between two DG–algebras.
-In addition, at the end of Section 3, we give several specific examples of graduations on group algebras and derivations that define the structure of the DG–algebra on them.
-The definitions and properties from combinatorial group theory that are necessary for the paper can be found in [LS01]
-1 Groupoid, characters and graded derivations.
-Let’s start with the definitions and fix the notation for future reference. We will study derivations over group algebras with a DG structure. A group algebra is defined in standard form as follows:
+In Section 1 we present the structure of a groupoid which is connected to the graded algebra structure.
+We will give a description of graded derivations as locally finite characters on this groupoid (see Theorem 1.1).
+In Section 2, we describe the inner, quasi-inner and central derivations in the graded case.
+Using this construction in the case of DG–algebras, in Section 3 we obtain a description of DG–algebra structures in terms of characters and an isomorphism criterion for DG–algebras.
+In addition, at the end of Section 3, we give several examples of gradings on group algebras and derivations that define a DG–algebra structure on them.
+The definitions and properties from combinatorial group theory that are necessary to understand the present paper can be found in [LS01]
+1 The groupoid, it’s characters, and graded derivations.
+Let’s start with the definitions and fix the notation for future reference. We will study derivations of group algebras with a DG structure. The standard definition of a group algebra is
 Definition 1.1.
-$K[G]$ – is the group algebra of a group $G$ over a field $K$ is an associative algebra over a field $K$, the elements of which are all possible formal finite sums of the form $\sum_{g\in G}k_{g}g,$ $g\in G,$ $k_{g}\in K$, and the operations are determined by the formulas:
+Let $G$ be a group and let $K$ a field.
+The group algebra $K[G]$ is the associative algebra over $K$ the elements of which are all possible formal finite sums of the form $\sum_{g\in G}k_{g}g,$ $g\in G,$ $k_{g}\in K$, and the operations are defined by:
 $\sum_{g\in G}a_{g}g+\sum_{g\in G}b_{g}g=\sum_{g\in G}(a_{g}+b_{g})g,$
-$(\sum_{g\in G}a_{g}g)(\sum_{g\in G}b_{g}g)=\sum_{h\in G}(\sum_{xy=h\in G}(a_{x}b_{y})h).$
-The elements of the group $G$ form the basis of the algebra $K[G]$. The multiplication of the basic elements in the group algebra is induced by the multiplication in group.
-Furthermore, we need the concept of a groupoid.
-Typically, this is a category in which every morphisms is invertible.
-More precisely, we will work with the following construction associated with a finitely generated group $G$, so this category is small (objects and morphisms are sets).
+$(\sum_{g\in G}a_{g}g)(\sum_{g\in G}b_{g}g)=\sum_{h\in G}\Big{(}\sum_{xy=h\in G}(a_{x}b_{y})h\Big{)}.$
+The elements of the group $G$ form a basis of $K[G]$. Multiplication of the basis elements in the group algebra is induced by the group multiplication.
+Furthermore, we need the notion of a groupoid.
+usually, it is defined as a category in which every morphism is invertible.
+We will work with the following construction associated with a finitely generated group $G$, so this category is small (the classes of objects and morphisms are sets).
 Definition 1.2.
-A groupoid $\Gamma$ contains a set of objects $Obj(\Gamma)$ and morphisms $Hom(\Gamma)$.
+A groupoid $\Gamma$ has a set of objects $\operatorname{Obj}(\Gamma)$ and a set of morphisms $\operatorname{Hom}(\Gamma)$.
 Each morphism has a source $s(\phi)$ and a target $t(\phi)$, i.e.
-each morphism is in the set $Hom(s(\phi),t(\phi))$.
+each morphism is in the set $\operatorname{Hom}(s(\phi),t(\phi))$.
 Composable morphisms have an associative operation (composition).
@@
   1. 1.
-    For each object $a\in Obj(\Gamma)$, there is a neutral endomorphism $1_{a}\in Hom(a,a)$ such that $\forall\phi\in Hom(b,a)$ $1_{a}\circ\phi=\phi$ and $\forall\phi\in Hom(a,b)$ $\phi\circ 1_{a}=\phi$.
+    For each object $a\in\operatorname{Obj}(\Gamma)$, there is a neutral endomorphism $1_{a}\in\operatorname{Hom}(a,a)$ such that $\forall\phi\in\operatorname{Hom}(b,a)$: $1_{a}\circ\phi=\phi$ and $\forall\phi\in\operatorname{Hom}(a,b)$ $\phi\circ 1_{a}=\phi$.
   2. 2.
-    There is an inverse for each morphism. I.e. $\forall\phi\in Hom(a,b)$ exists $\psi\in Hom(b,a)$ such that $\phi\circ\psi=1_{b}$ and $\psi\circ\phi=1_{a}$.
+    Each morphism has the inverse: fir every $\phi\in\operatorname{Hom}(a,b)$ there exists $\psi\in\operatorname{Hom}(b,a)$ such that $\phi\circ\psi=1_{b}$ and $\psi\circ\phi=1_{a}$.
 1.1 Groupoid of conjugacy action.
-The following construction is a generalization of the structure from [AMS17].
-Suppose we have a graded group algebra $\mathbb{C}[G]\cong\bigoplus A_{i}$, here $i\in\mathbb{Z}$.
-Group $G$ assumes to be finitely generated.
-We will define the groupoid of conjugacy action, denoted by $\Gamma$, for the group $G$ in the sense of Definition 1.2.
+The following construction generalizes the one from [AMS17].
+Suppose we have a $\mathbb{C}$–graded group algebra $\mathbb{C}[G]\cong\bigoplus A_{i}$.
+The group $G$ is assume to be finitely generated.
+We will define the groupoid of conjugacy action $\Gamma$ for the group $G$ (in the sense of Definition 1.2) as follows:
 […]
-Consider two morphisms $\phi=(u_{1},v_{1})$ and $\psi=(u_{2},v_{2})$ such that $t(\phi)=s(\psi)$, that is, the composition is defined for them $\psi\circ\phi$.
-Define the composition of morphisms as follows
-$(u_{2},v_{2})\circ(u_{1},v_{1}):=((-1)^{|v_{2}|}v_{2}u_{1},v_{2}v_{1})).$
-This formula is conveniently recorded in the form of a diagram.
+Consider two morphisms $\phi=(u_{1},v_{1})$ and $\psi=(u_{2},v_{2})$ such that $t(\phi)=s(\psi)$, that is, the composition $\psi\circ\phi$ should be defined. Define it as follows:
+$(u_{2},v_{2})\circ(u_{1},v_{1}):=((-1)^{|v_{2}|}v_{2}u_{1},v_{2}v_{1}).$
+This formula is conveniently depicted as a diagram.
 $(-1)^{|v_{1}|}v_{1}^{-1}u_{1}=u_{2}v_{2}^{-1}$$u_{1}v_{1}^{-1}$$(-1)^{|v_{2}|}v_{2}^{-1}u_{2}$$\scriptstyle{(u_{2},v_{2})}$$\scriptstyle{(u_{1},v_{1})}$$\scriptstyle{(u_{2},v_{2})\circ(u_{1},v_{1})}$
@@
 First of all, we note that $\exists g\rightarrow h\Leftrightarrow\exists t:h=(-1)^{|t|}tgt^{-1}.$
-Let we have $\phi=(u_{1},v_{1})$ and $\psi=(u_{2},v_{2})$ such that $t(\phi)=s(\psi)$. Let’s say $s(\phi)=a$, $t(\phi)=s(\psi)=b$ and $t(\psi)=c$.
-Check that
+Consider $\phi=(u_{1},v_{1})$ and $\psi=(u_{2},v_{2})$ such that $t(\phi)=s(\psi)$. Define $s(\phi)=a$, $t(\phi)=s(\psi)=b$ and $t(\psi)=c$.
+We need to check that
 $s(\psi\circ\phi)=a$ and $t(\psi\circ\phi)=c$.
-By definition
-$s(\psi\circ\phi)=(-1)^{|v_{2}v_{1}|}v_{1}^{-1}v_{2}^{-1}(-1)^{|v_{2}|}v_{2}u_{1}$ = $(-1)^{|v_{1}|}v_{1}^{-1}u_{1}=a$.
+By definition,
+$s(\psi\circ\phi)=(-1)^{|v_{2}v_{1}|\mathstrut}v_{1}^{-1\mathstrut}v_{2}^{-1\mathstrut}(-1)^{|v_{2}|\mathstrut}v_{2}u_{1}$ = $(-1)^{|v_{1}|\mathstrut}v_{1}^{-1\mathstrut}u_{1}=a$.
 Calculate that
-$t(\psi\circ\phi)$ = $(-1)^{|v_{2}|}v_{2}u_{1}v_{1}^{-1}v_{2}^{-1}=u_{2}v_{2}^{-1}=c$, because $u_{1}v_{1}^{-1}=(-1)^{|v_{2}|}v_{2}^{-1}u_{2}$
-Now we prove the associativity of the composition. Let $\phi_{1},\phi_{2},\phi_{3}$ - morphisms such that $s(\phi_{1})=t(\phi_{2})$ and $s(\phi_{2})=t(\phi_{3})$, then
+$t(\psi\circ\phi)$ = $(-1)^{|v_{2}|\mathstrut}v_{2}u_{1}v_{1}^{-1\mathstrut}v_{2}^{-1\mathstrut}=u_{2}v_{2}^{-1\mathstrut}=c$, because $u_{1}v_{1}^{-1\mathstrut}=(-1)^{|v_{2}|\mathstrut}v_{2}^{-1\mathstrut}u_{2}$
+Here we shall prove that the composition is associative. Let $\phi_{1},\phi_{2},\phi_{3}$ be morphisms such that $s(\phi_{1})=t(\phi_{2})$ and $s(\phi_{2})=t(\phi_{3})$. We need to show that
 $\phi_{1}\circ(\phi_{2}\circ\phi_{3})=(\phi_{1}\circ\phi_{2})\circ\phi_{3}.$
-We proved above that all compositions are defined correctly. Let’s calculate the left part.
-$\phi_{2}\circ\phi_{3}=((-1)^{|v_{2}|}v_{2}u_{3},v_{2}v_{3}))$ $\phi_{1}\circ(\phi_{2}\circ\phi_{3})=(u_{1},v_{1})\circ((-1)^{|v_{2}|}v_{2}u_{3},v_{2}v_{3}))=((-1)^{|v_{1}|+|v_{2}|}v_{1}v_{2}u_{3},v_{1}v_{2}v_{3}).$
-Now let’s calculate the right side of the expression
-$(\phi_{1}\circ\phi_{2})\circ\phi_{3}=((-1)^{|v_{1}|}v_{1}u_{2},v_{1}v_{2})\circ(u_{3},v_{3})=((-1)^{|v_{1}|+|v_{2}|}v_{1}v_{2}u_{3},v_{1}v_{2}v_{3}).$
-The left and right sides of the expression are equal, which means that the operation $\circ$ is associative.
-Let’s show that there is an element $1_{g}$, such that:
-$s(1_{g})=t(1_{g})=g$
-$\phi\cdot 1_{s(\phi)}=\phi=1_{t(\phi)}\cdot\phi$, $\forall\phi\in$ $Hom(\Gamma)$
-Let’s take as a $1_{g}$ element $(g,e)$, where e is a trivial element of the group G.
+We have proved above that these compositions are well–defined. The left side is
+$\phi_{2}\circ\phi_{3}=((-1)^{|v_{2}|}v_{2}u_{3},v_{2}v_{3}))$ $\phi_{1}\circ(\phi_{2}\circ\phi_{3})=(u_{1},v_{1})\circ((-1)^{|v_{2}|}v_{2}u_{3},v_{2}v_{3}))=((-1)^{|v_{1}|+|v_{2}|}v_{1}v_{2}u_{3},v_{1}v_{2}v_{3}),$
+While the right one is equal to it,
+$(\phi_{1}\circ\phi_{2})\circ\phi_{3}=((-1)^{|v_{1}|}v_{1}u_{2},v_{1}v_{2})\circ(u_{3},v_{3})=((-1)^{|v_{1}|+|v_{2}|}v_{1}v_{2}u_{3},v_{1}v_{2}v_{3}),$
+proving associativity.
+Let us show that there is an element $1_{g}$ such that
+$s(1_{g})=t(1_{g})=g$,
+$\phi\cdot 1_{s(\phi)}=\phi=1_{t(\phi)}\cdot\phi$, $\forall\phi\in$ $\operatorname{Hom}(\Gamma).$
+Take $1_{g}=(g,e)$, where $e$ is the identity of the group $G$.
 $(-1)^{|e|}e^{-1}g=g=ge^{-1}$$\scriptstyle{(g,e)}$
-Indeed, let’s take an arbitrary morphism $(u,v)$, than:
-$(uv^{-1},e)(u,v)=(u,v)$.
+Indeed, choosing an arbitrary morphism $(u,v)$, we have:
+$(uv^{-1},e)(u,v)=(u,v)$;
 $(u,v)((-1)^{|v|}v^{-1}u,e)=((-1)^{|v|}v(-1)^{|v|}v^{-1}u,e)=(u,v)$.
-We have shown that $\Gamma$ is a category. Now it is necessary to check the inversibility of the arrows.
-We show the presence of an inverse morphism $\forall\phi\in Hom(\Gamma)$. For morphism of the form (u,v) let’s call the inverse morphism $((-1)^{|v|}v^{-1}uv^{-1},v^{-1})$.
-Indeed:
-$(u,v)((-1)^{|v|}v^{-1}uv^{-1},v^{-1})=(uv^{-1},e)=(t(\phi),e).$
+We have now shown that $\Gamma$ is a category. It remains to check the invertibility of the arrows.
+We claim the the inverse of a morphism of the form $(u,v)$ is $((-1)^{|v|}v^{-1}uv^{-1},v^{-1})$.
+Indeed,
+$(u,v)((-1)^{|v|}v^{-1}uv^{-1},v^{-1})=(uv^{-1},e);=(t(\phi),e).$
 $((-1)^{|v|}v^{-1}uv^{-1},v^{-1})(u,v)=((-1)^{|v|}v^{-1}u,e)=(s(\phi),e).$
 $(-1)^{|v|}v^{-1}u$$uv^{-1}$$\scriptstyle{(u,v)}$$\scriptstyle{((-1)^{|v|}v^{-1}uv^{-1},v^{-1})}$
-Thus, the Proposition 1.1. is proved. ∎
+This proves Proposition 1.1. ∎
 The following properties of the groupoid are easily verified by direct calculation.
 Proposition 1.2.
-For a groupoid $\Gamma$ holds:
   1. 1.
-    For endomorphisms $Hom(a,a),a\in\pm G$ holds the following descriptions
-    $Hom(a,a)=\{\chi\in Hom(\Gamma)\mid s(\chi)=(-1)^{|t|}t^{-1}a,$ $t(\chi)=at^{-1},$ $t\in Z(a)$ and $t\in$ $A_{2i}[G]\}$
+    The endomorphisms $\operatorname{Hom}(a,a),a\in\pm G$, admit the following description:
+    $\operatorname{Hom}(a,a)=\{\chi\in\operatorname{Hom}(\Gamma)\mid s(\chi)=(-1)^{|t|}t^{-1}a,$ $t(\chi)=at^{-1},$ $t\in Z(a)$ and $t\in$ $A_{2i}[G]\}$
   2. 2.
-    For $a]:=\{b\in\pm G\mid\exists t\in\pm G:b=(-1)^{|t|}t^{-1}at\}$ denote by $\Gamma_{[a]}$, a subgroupoid such that $Obj(\Gamma_{[a]})=[a]$. The following properties holds:
+    For $a]:=\{b\in\pm G\mid\exists t\in\pm G:b=(-1)^{|t|}t^{-1}at\}$, denote by $\Gamma_{[a]}$ the subgroupoid such that $\operatorname{Obj}(\Gamma_{[a]})=[a]$. The following properties hold:
     1. (a)
-      $Hom(\Gamma_{[a]})=\{\phi\in Hom(\Gamma)\mid s(\phi)\in[a],t(\phi)\in[a]\}$.
+      $\operatorname{Hom}(\Gamma_{[a]})=\{\phi\in\operatorname{Hom}(\Gamma)\mid s(\phi)\in[a],t(\phi)\in[a]\}$.
     2. (b)
@@
   3. 3.
-    There is a left action $Hom(a,a)\times Hom(a,b)\rightarrow Hom(a,b)$, $(\phi,\psi)\to\psi\circ\phi$. This action is free and transitive.
-Let us show that with the help of characters of $\Gamma$, it is possible to describe a derivation satisfying the graded Leibniz rule.
+    The left action $\operatorname{Hom}(a,a)\times\operatorname{Hom}(a,b)\rightarrow\operatorname{Hom}(a,b)$: $(\phi,\psi)\to\psi\circ\phi$ is free and transitive.
+Let us show that, with the help of characters of $\Gamma$, it is possible to give a description of derivations satisfying the graded Leibniz rule.
 Definition 1.3.
-We say that the linear operator $d$: $\mathbb{C}[G]$ $\to$ $\mathbb{C}[G]$ satisfies the graded Leibniz rule if
+The linear operator $d$: $\mathbb{C}[G]$ $\to$ $\mathbb{C}[G]$ satisfies the graded Leibniz rule if
 $d(uv)=d(u)v+(-1)^{|u|}ud(v)$, $\forall u,v\in A_{i}$
 Definition 1.4.
-We call the map $\chi$ : $Hom(\Gamma)$ $\to$ $\mathbb{C}$ a character if it satisfies the identity $\chi(\psi\circ\phi)=\chi(\psi)+\chi(\phi)$
+The map $\chi$ : $\operatorname{Hom}(\Gamma)$ $\to$ $\mathbb{C}$ is a character if it satisfies the identity $\chi(\psi\circ\phi)=\chi(\psi)+\chi(\phi)$:
 $\bullet$$\bullet$$\bullet$$\scriptstyle{\chi(\psi)}$$\scriptstyle{\chi(\phi)}$$\scriptstyle{\chi(\psi\circ\phi)=\chi(\psi)+\chi(\phi)}$
 Definition 1.5.
-We will say that the character $\chi$ is locally finite if for all $v\in pmG$ we get that $\chi(u,v)=0$ for all except a finite number $u$.
-We will denote by $\chi(\Gamma)$ the space of local finite characters. In section 1.2 we will show that $\chi(\Gamma)$ has a natural structure of Lie algebra.
-The graded Leibniz rule in a group algebra can be written in terms of characters on the groupoid $\Gamma$ as follows.
+A character $\chi$ is locally finite if for all $v\in\pm G$ we have $\chi(u,v)=0$ for all except a finite number of $u$.
+We will denote by $\chi(\Gamma)$ the space of locally finite characters. In section 1.2 we will show that $\chi(\Gamma)$ has a natural Lie algebra structure.
+The graded Leibniz rule in a group algebra can be expressed in terms of characters on the groupoid $\Gamma$ as follows.
 Proposition 1.3.
-For any graded derivation $d$ there exists a locally finite character such that
-$d(u)=\sum\limits_{g\in G}\Big{(}\sum\limits_{h\in\pm G}\chi(\phi(h,g)\lambda^{h}\Big{)}g,\forall u\in A_{i}$ (1)
+For any graded derivation $d$ there exists a locally finite character $\chi$ such that
+$d(u)=\sum\limits_{g\in G}\Big{(}\sum\limits_{h\in\pm G}\chi(\phi(h,g)\lambda^{h}\Big{)}g;\forall u\in A_{i}.$ (1)
 Proof.
-It is enough to check the statement on the elements $g\in\pm G$, since they generate $\mathbb{C}[G]$, as a vector space.
-For any element $g$ of group G we have $d(g)=\sum\limits_{h\in G}d_{g}^{h}h$
-Define $\chi(\phi(h,g)):=d_{g}^{h}$, $\chi(\phi(-h,g)):=-d_{g}^{h}$ and let us show that it is a character. We have
-$d(g_{2}g_{1})=d(g_{2})g_{1}+(-1)^{|g_{2}|}g_{2}d(g_{1}).$
-Let’s rewrite this in a new notation:
+It is enough to check the statement on the elements $g\in\pm G$, since they generate $\mathbb{C}[G]$ as a vector space.
+For any element $g$ of group G we have $d(g)=\sum\limits_{h\in G}d_{g}^{h}h$.
+Let us define $\chi(\phi(h,g)):=d_{g}^{h}$, $\chi(\phi(-h,g)):=-d_{g}^{h}$ and show that $\chi$ it is a character. We have
+$d(g_{2}g_{1})=d(g_{2})g_{1}+(-1)^{|g_{2}|}g_{2}d(g_{1}),$
+or after a basis expansion
 $\sum\limits_{h\in G}d_{g_{2}g_{1}}^{h}h=\sum\limits_{h\in G}d_{g_{2}}^{h}hg_{1}+(-1)^{|g_{2}|}g_{2}\sum\limits_{h\in G}d_{g_{1}}^{h}h.$
-It is clear that summation over all $h$ from the group for $h$ and $gh$ is equivalent, so we can convert the right-hand sides to the following form:
+It is clear that summing over all $h$ and all $gh$ is equivalent, so we can convert the right-hand side to the following form:
 $\sum\limits_{h\in G}d_{g_{2}g_{1}}^{h}h=\sum\limits_{h\in G}d_{g_{2}}^{hg_{1}^{-1}}h+\sum\limits_{h\in G}d_{g_{1}}^{(-1)^{|g_{2}|}g_{2}^{-1}h}h\Rightarrow d_{g_{2}g_{1}}^{h}=d_{g_{2}}^{hg_{1}^{-1}}+d_{g_{1}}^{(-1)^{|g_{2}|}g_{2}^{-1}h}.$ (2)
-Let’s make a replacement:
+Let’s make a substitution:
 $h_{1}=hg_{1}^{-1},$
@@
 It remains only to note that $((-1)^{|g_{2}|}g_{2}h_{2},g_{2}g_{1})=(h_{1},g_{2})(h_{2},g_{1}).$
-A corollary of the above statement is the following theorem, the proof of which is similar to the non-graded case, see Theorem 1 of [Aru20].
+A corollary of the above statement is the following theorem, proved similarly to the non-graded case, see Theorem 1 of [Aru20].
 Theorem 1.1.
-For any graded derivation $d$ there exists such a unique locally finite character of $\chi_{d}$, that the formula is valid
+For any graded derivation $d$ there exists a unique locally finite character $\chi_{d}$ satisfying
 $d(a)=(-1)^{|a|}a\Big{(}\sum\limits_{t\in G}\chi((-1)^{|a|}at,a)t\Big{)}.$
-In fact, we can say that we have an isomorphism $Der(\mathbb{C}[G])\cong\chi(\Gamma)$.
-To prove this, we need to use formula (1) to element $a\in\pm G$ and write an arbitrary element $g\in G$ as $g=at$ and then remove the terms with a zero coefficient from the sum.
+There is therefore an isomorphism $\operatorname{Der}(\mathbb{C}[G])\cong\chi(\Gamma)$.
+Proof.
+To prove this, we need to apply formula (1) to $a\in\pm G$, write an arbitrary element $g\in G$ as $g=at$ and then drop the terms with a zero coefficient from the sum.
 See the proof of Theorem 1 in [Aru20] for more details.
+∎
 1.2 Characters.
-Locally finite characters are isomorphic to derivations, which means that the binary commutator operation on derivations generates a binary operation on characters.
-Formula (1) allows us to define a composition in character space.
-Let $d_{1},d_{2}$ be the derivations, defined by the characters $\chi_{d_{1}},\chi_{d_{2}}$.
-The character that corresponds to the commutator of derivations $d_{1},d_{2}]$ is denoted by $\chi_{[d_{1},d_{2}]}$.
-It defines an operation on the character space.
-$\{\chi_{d_{1}},\chi_{d_{2}}\}:=\chi_{[d_{1},d_{2}]}$ (3)
-For this operation, we have the following equality
+Locally finite characters are isomorphic to derivations, which means that the commutator of derivations yields a binary operation on characters.
+Formula (1) allows us to define a commutator in character space.
+Let $d_{1},d_{2}$ be the derivations defined by the characters $\chi_{d_{1}},\chi_{d_{2}}$.
+Denote the character that corresponding to the commutator $d_{1},d_{2}]$ by $\chi_{[d_{1},d_{2}]}$.
+This defines an operation on the space of characters.
+$\{\chi_{d_{1}},\chi_{d_{2}}\}:=\chi_{[d_{1},d_{2}]}.$ (3)
+For this operation, we have the following equality:
 Proposition 1.1.
-Holds the formula
+The operation on characters above can be expressed as
 $\{\chi_{d_{1}},\chi_{d_{2}}\}(a,g)=\sum\limits_{h\in G}\chi_{d_{1}}(a,h)\chi_{d_{2}}(h,g)-\chi_{d_{2}}(a,h)\chi_{d_{1}}(h,g).$ (4)
 Proof.
-Let $g\in G$, than we have
+Let $g\in G$. Then we have
 $d_{1}(g)$ $=\sum\limits_{h\in G}\chi_{d_{1}}(h,g)h,$
@@
 $d_{2}d_{1}(g)=\sum\limits_{h\in G}\chi_{d_{1}}(h,g)\Big{(}\sum\limits_{a\in\pm G}\chi_{d_{2}}(a,h)a\Big{)}$.
-Due to local finiteness we can change the order of summation in the last equation and get
+Thanks to local finiteness, we can swap the sums in the last equation and get
 $d_{1},d_{2}](h)=\sum\limits_{a\in G}\Big{(}\sum\limits_{h\in G}\chi_{d_{2}}(h,g)\chi_{d_{1}}(a,h)-\chi_{d_{1}}(h,g)\chi_{d_{2}}(a,h)\Big{)}a$.
-Equation $\{\chi_{d_{1}},\chi_{d_{2}}\}(a,g)$ it is coefficient under $a$, so
+The value of $\{\chi_{d_{1}},\chi_{d_{2}}\}(a,g)$ is the coefficient of $a$, so
 $\{\chi_{d_{1}},\chi_{d_{2}}\}(a,g)=\sum\limits_{h\in G}\chi_{d_{1}}(a,h)\chi_{d_{2}}(h,g)-\chi_{d_{2}}(a,h)\chi_{d_{1}}(h,g)$.
-It follows that the pair $(\chi_{\Gamma},\{\cdot;\cdot\})$ is a Lie algebra which is isomorphic by proposition 1.3 to the algebra of all graded derivations.
-2 Examples of graded derivations and their description using characters on a groupoid.
-We will demonstrate the correspondence between characters on groupoid and graded derivations using examples of inner derivations, central derivations and quasi-inner derivations.
-It gives give easy examples of such operators.
+It follows that the $(\chi_{\Gamma},\{\cdot;\cdot\})$ is a Lie algebra, isomorphic to the algebra of all graded derivations by proposition 1.3.
+2 Examples of graded derivations and their description in terms of characters.
+We will illustrate the correspondence between characters on a groupoid and graded derivations using examples of inner derivations, central derivations and quasi-inner derivations.
+They will serve as easy examples of such operators.
 2.1 Inner derivations
-Consider the element $a\in\pm G$. Let’s define the mapping $\chi^{a}:Hom(\Gamma)\rightarrow\mathbb{C}$ as follows (assuming that $a\neq b$):
-$\chi^{a}(\phi):=\begin{cases}1,&\text{$\phi\in Hom(b,a)$}\\ -1,&\text{$\phi\in Hom(a,b)$}\\ 0,&\text{else.}\end{cases}$ (5)
-It is clear that $\chi^{a}$ is trivial on endomorphisms.
+Consider an element $a\in\pm G$. Define the mapping $\chi^{a}:\operatorname{Hom}(\Gamma)\rightarrow\mathbb{C}$ as follows (assuming $a\neq b$):
+$\chi^{a}(\phi):=\begin{cases}1,&\text{$\phi\in\operatorname{Hom}(b,a)$}\\ -1,&\text{$\phi\in\operatorname{Hom}(a,b)$}\\ 0,&\text{otherwise.}\end{cases}$ (5)
+Clearly $\chi^{a}$ is trivial on endomorphisms.
 Proposition 2.1.
-Map $\chi^{a}$ sets the inner derivations for $a\in\pm G$ by formula
-$d_{a}:x\rightarrow[a,x]_{grad}$ , where
-$a,x]_{grad}=ax-(-1)^{|x|}xa$.
+The map $\chi^{a}$ defines an inner derivations for $a\in\pm G$ by the formula
+$d_{a}:x\mapsto[a,x]_{grad}=ax-(-1)^{|x|}xa$.
 Proof.
@@
 We will now give an example of a derivation that is not inner.
-It will be demonstrated that such a derivation has an easy calculation form and is not trivial on endomorphisms.
-It will thus be established that they cannot be presented as a (perhaps formal) sum of inner derivations.
+We will show that it admits an simple explicit form and is not trivial on endomorphisms.
+That entails that it cannot be expressed as a (perhaps formal) sum of inner derivations.
 Definition 2.1.
-We say that map $\tau:G\rightarrow\mathbb{C}$ – is graded group character, if it satisfies the following property:
+That the map $\tau:G\rightarrow\mathbb{C}$ is a graded group character if it satisfies the following property:
 $\tau(ab)=\tau(a)+(-1)^{|a|}\tau(b),$ $\forall a,b\in G$.
-Let’s fix the graded group character $\tau$, and let $z\in Z(G)$ be an element of group center.
+Let’s fix a graded group character $\tau$, and let $z\in Z(G)$ be central element of the group.
 Definition 2.2.
-Define map $d_{z}^{\tau}:\mathbb{C}[G]\rightarrow\mathbb{C}[G]$ on the generators $g\in\pm G$ as
-$d_{z}^{\tau}:g\rightarrow\tau(g)gz.$
-Continued by linearity for the whole algebra $\mathbb{C}[G]$ map $d_{z}^{\tau}$ we will call it as the central graded derivation.
+Define the map $d_{z}^{\tau}:\mathbb{C}[G]\rightarrow\mathbb{C}[G]$ on the generators $g\in\pm G$ as
+$d_{z}^{\tau}:g\rightarrow\tau(g)gz$
+and extend it by linearity to the whole algebra $\mathbb{C}[G]$. We will call it a central graded derivation.
 Proposition 2.2.
@@
 Proof.
-[…]
+It is enough to check the graded Leibniz rule on arbitrary generators $u,v\in G$:
 $d_{z}^{\tau}(uv)=\tau(uv)uvz=\tau(u)uzv+(-1)^{|u|}u\tau(v)vz=d_{z}^{\tau}(u)v+(-1)^{|u|}ud_{z}^{\tau}(v)$.
-On arbitrary elements of a group algebra, the graded Leibniz rule is fulfilled due to the linearity of the $d_{z}^{\tau}$. ∎
+On arbitrary elements of a group algebra, the graded Leibniz rule is satisfied due to the linearity of $d_{z}^{\tau}$. ∎
 Proposition 2.3.
@@
 Proof.
-Let $\chi_{\tau,z}$ – a character that corresponds to derivation $d_{z}^{\tau}$. By corollary 1 we have equality
+Let $\chi_{\tau,z}$ be a character that corresponds to the derivation $d_{z}^{\tau}$. By corollary 1 we have the equality
 $d_{z}^{\tau}(g)=\tau(g)gz=g\Big{(}\sum\limits_{t\in G}\chi((-1)^{|g|}gt,g)(-1)^{|g|}t\Big{)}$.
-Therefore, the character is not equal to 0 only in the case of $t=z$:
+Therefore, this character is not equal to 0 only in the case $t=z$:
 $\chi((-1)^{|g|}gz,g)=(-1)^{|g|}\tau(g)$.
-Denote $\phi=((-1)^{|g|}gz,g)$ and calculate its beginning and end
+Define the morphism $\phi=((-1)^{|g|}gz,g)$ and calculate its beginning and end
 $s(\phi)=z$,
 $t(\phi)=(-1)^{|g|}z.$
-Then in the case when $g\in A_{2n}$, we have: $s(\phi)=t(\phi)$ $\Rightarrow\chi_{\tau,z}$ non trivial on a loop $\phi$, and $d_{z}^{\tau}$ it is not an inner derivation. ∎
-A direct check shows that inner and central derivations are subalgebras. For more details see [Aru20].
+Then in the case when $g\in A_{2n}$, we have: $s(\phi)=t(\phi)$ $\Rightarrow\chi_{\tau,z}$ is non–trivial on a loop $\phi$, thus $d_{z}^{\tau}$ it is not an inner derivation. ∎
+A direct check shows that inner and central derivations constitute subalgebras. For more details see [Aru20].
 2.3 Quasi-inner derivations
-As seen from formula (5) inner derivations are trivial on endomorphisms; however, the opposite is not true. We will call the class of such derivations quasi-inner.
+As seen from the formula (5), inner derivations are trivial on endomorphisms; however, the converse is not true. We will call such derivations quasi-inner.
 Definition 2.3.
-Define QInnDer($\mathbb{C}[G]$) derivation as follows:
-$QInnDer(\mathbb{C}[G])=\left\{d\in Der\mid\forall a\in Obj(\Gamma),\forall\phi\in Hom(a,a)\Rightarrow\chi_{d}(\phi)=0\right\}.$ (6)
-Quasi-inner derivations are clearly a sub-space, but the more significant statement is that they constitute an ideal in algebra of all derivations.
+Define $\operatorname{QInnDer}(\mathbb{C}[G])$ as follows:
+$\operatorname{QInnDer}(\mathbb{C}[G])=\left\{d\in\operatorname{Der}\mid\forall a\in\operatorname{Obj}(\Gamma),\forall\phi\in\operatorname{Hom}(a,a):\chi_{d}(\phi)=0\right\}.$ (6)
+Quasi-inner derivations are clearly a subspace, but the more significant statement is that they constitute an ideal in the algebra of all derivations.
 Proposition 2.4.
-QInnDer($\mathbb{C}[G]$) $\triangleleft$ Der($\mathbb{C}[G]$) is an ideal.
+$\operatorname{QInnDer}(\mathbb{C}[G])$ $\triangleleft$ $\operatorname{Der}(\mathbb{C}[G])$ is an ideal.
 Proof.
-Let’s proof that QInnDer($\mathbb{C}[G]$) is a subalgebra, i.e.
-$d_{1}$, $d_{2}$ $\in$ QInnDer($\mathbb{C}[G]$) $\Rightarrow$ $\left[d_{1},d_{2}\right]\in QInnDer(\mathbb{C}[G])$.
-character $\chi_{d_{1}}$ can be represented as:
+Let us prove that $\operatorname{QInnDer}(\mathbb{C}[G])$ is a subalgebra, i.e.
+$d_{1}$, $d_{2}$ $\in$ $\operatorname{QInnDer}(\mathbb{C}[G]$) $\Rightarrow$ $\left[d_{1},d_{2}\right]\in\operatorname{QInnDer}(\mathbb{C}[G])$.
+The character $\chi_{d_{1}}$ can be represented as
 $\chi_{d_{1}}=\sum\limits_{a\in G}\lambda^{a}\chi^{a}$, $\lambda^{a}\in\mathbb{C}$,
-since $\chi_{d_{1}}$ is trivial on endomorphisms, where $\chi_{d_{1}}$ defined by formula (5). Similarly with a character $\chi_{d_{2}}$:
+where $\chi_{d_{1}}$ defined by formula (5), since $\chi^{a}$ is trivial on endomorphisms. Similarly for the character $\chi_{d_{2}}$:
 $\chi_{d_{2}}=\sum\limits_{b\in G}\lambda^{b}\chi^{b}$, $\lambda^{b}\in\mathbb{C}$.
@@
 $\{\chi^{a},\chi^{b}\}=\chi^{ab}-\chi^{ba}$.
-Obtain the final formula for $\{\chi_{d_{1}},\chi_{d_{2}}\}$:
+Claim:
 $\{\chi_{d_{1}},\chi_{d_{2}}\}=\sum\limits_{a\in G}\sum\limits_{b\in G}\lambda^{a}\mu^{b}\chi^{ab}-\sum\limits_{a\in G}\sum\limits_{b\in G}\lambda^{a}\mu^{b}\chi^{ba}$,
-$\{\chi_{d_{1}},\chi_{d_{2}}\}\in QInnDer(\mathbb{C}[G])$, to proof that, consider the value of the character on the loop $(uz,z),z\in Z_{G}(u)$
+thus $\{\chi_{d_{1}},\chi_{d_{2}}\}\in\operatorname{QInnDer}(\mathbb{C}[G])$. To prove that, consider the value of the character on the loop $(uz,z),z\in Z_{G}(u)$:
 $\{\chi_{d_{1}},\chi_{d_{2}}\}(uz,z)=\sum\limits_{ab=zuz^{-1}}\lambda^{a}\mu^{b}-\sum\limits_{ab=u}\lambda^{a}\mu^{b}-\sum\limits_{ba=u}\lambda^{a}\mu^{b}+\sum\limits_{ba=zuz^{-1}}\lambda^{a}\mu^{b}=0$.
-Now pass directly to the proof of the theorem. Represent $\chi_{d_{0}}$ as
+We will now turn to the proof of the theorem. Represent $\chi_{d_{0}}$ as
 $\chi_{d_{0}}=\sum\limits_{a\in G}\lambda^{a}\chi^{a}$.
-Prove the theorem for $\chi^{a}$ and extend the result to the $\chi_{d_{0}}$, using the bilinearity of the commutator.
-Consider character $\{\chi_{d},\chi^{a}\}$.
-Proof that it is trivial on endomorphisms, i.e.
-$\forall b\in G$ and for $\forall z\in Z_{G}(b)$ is performed as $\{\chi_{d},\chi^{a}\}(bz,z)=0$.
+It is sufficient to prove the theorem for $\chi^{a}$ and extend the result to the $\chi_{d_{0}}$ using bilinearity of the commutator.
+Consider the character $\{\chi_{d},\chi^{a}\}$.
+We need to prove that it is trivial on endomorphisms, i.e.
+that for all $b\in G$ and for $z\in Z_{G}(b)$ it satisfies $\{\chi_{d},\chi^{a}\}(bz,z)=0$.
 By Proposition 3.1:
 $\{\chi_{d},\chi^{a}\}(bz,z)=\sum\limits_{h\in G}\chi_{d}(bz,h)\chi^{a}(h,z)-\chi^{a}(bz,h)\chi_{d}(h,z)$.
-$\chi^{a}(h,z)\neq 0$ only in 2 cases: when $h=za$ and $h=az$. $\chi^{a}(bz,h)\neq 0$ only in 2 cases: when $h=bza^{-1}$ and $h=a^{-1}bz$. It means that
+Here $\chi^{a}(h,z)\neq 0$ only in 2 cases: when $h=za$ and when $h=az$; and $\chi^{a}(bz,h)\neq 0$ only in 2 cases: when $h=bza^{-1}$ and when $h=a^{-1}bz$. This means that
 $\{\chi_{d},\chi^{a}\}(bz,z)=\chi_{d}(bz,za)-\chi_{d}(bz,az)+\chi_{d}(a^{-1}bz,z)-\chi_{d}(bza^{-1},z)$.
@@
 $(a^{-1}bz,z)\circ(bz,za)=(bz,az)\circ(bza^{-1},z)$.
-That means
+Hence
 $\chi_{d}(bz,za)+\chi_{d}(a^{-1}bz,z)=\chi_{d}(bz,az)+\chi_{d}(bza^{-1},z)$,
 $\{\chi_{d},\chi^{a}\}(bz,z)=0$.
-It is easy to see that central derivations are not quasi-inner, for which it is sufficient to repeat the reasoning from proposition 2.2.
+It is easy to see that central derivations are not quasi-inner by repeating the reasoning from proposition 2.3.
 3 Description of the structure of a DG–algebra in terms of characters on a groupoid.
-Our next goal is to give the description of possible DG–algebra sctructure up to isomorphism. We follow the definitions from [FHT01]
+Our next goal is to give a description of the possible DG–algebra structures up to isomorphism. We follow the definitions of [FHT01].
 Definition 3.1.
-Differential graded algebra (DG–algebra) is a graded algebra $A\cong\bigoplus A_{i}$, equipped with derivation
-$d:A\rightarrow A$,
-which has a degree of $1$ (the case of a cochain complex) or 1-1 (the case of a chain complex) and satisfies following properties
+Let $A\cong\bigoplus A_{i}$ be a $\mathbb{Z}$–graded algebra equipped with a derivation
+$d:A\rightarrow A$
+of degree $1$ (the case of a cochain complex) or 1-1 (the case of a chain complex) and satisfying the following conditions:
   1. 1.
@@
     $d(uv)=d(u)v+(-1)^{|u|}ud(v)$.
-We will give the necessary and sufficient conditions from the point of view of characters on a groupoid that derivatiation sets the structure of a DG–algebra on a graded group algebra.
-We will describe the case of a cochain complex, since the reasoning is similar for a chain complex.
+We say that $(A,d)$ is a differential graded algebra or DG–algebra.
+We will give the necessary and sufficient conditions from the point of view of groupoid characters for a derivation to define a DG–algebra structure on a graded group algebra.
+We will describe the cochain complex case, since the reasoning in the chain complex is similar.
 Theorem 3.1.
-The characters on the groupoid define the structures of the DG–algebra if and only if the following conditions are satisfied:
+A character $\chi$ on the groupoid $\Gamma$ defines a DG–algebra structure if and only if the following conditions are satisfied:
   1. 1.
@@
     For all $(h,g)\in Hom(\Gamma_{A_{i}})$, $i\neq 1$, we have $\chi(h,g)=0$.
-The main idea is to prove that the first condition is equivalent to $d^{2}(g)=0$, and that the second condition is equivalent to $d(A^{i})\subseteq A^{i+1}$.
+The main idea is to prove that the first condition is equivalent to $d^{2}(g)=0$ and the second one is equivalent to $d(A^{i})\subseteq A^{i+1}$.
 Proof.
-Using proposition 2.2, we rewrite the properties of the DG–algebra in terms of characters.
+Using proposition 2.2, we rewrite the defining properties of a DG–algebra in terms of characters.
   1. 1.
@@
     $d^{2}(g)=d\Big{(}\sum\limits_{h\in G}d^{h}_{g}h\Big{)}=\sum\limits_{h\in G}d^{h}_{g}\Big{(}\sum\limits_{x\in G}d^{x}_{h}x\Big{)}=\sum\limits_{h\in G}(h,g)\Big{(}\sum\limits_{x\in G}(x,h)x\Big{)}=\sum\limits_{x\in G}\Big{(}\sum\limits_{h\in G}(\chi(h,g)\chi(x,h))\Big{)}x.$
-    So the condition (7) holds.
   2. 2.
-    Let $g\in A^{i}$, than $d(g)=\sum\limits_{h\in G}(h,g)h\in A^{i+1}\Leftrightarrow(h,g)=0$, if $h\notin A^{i+1}$.
-    Note that if $u\in A^{i+1},v\in A^{i}$, than for $\phi=(u,v)$ we have: $s(\phi)=(-1)^{|v|}v^{-1}u\in A^{1},t(\phi)=uv^{-1}\in A^{1}$,this means that in the case of a DG algebra, the characters may not be equal to 0 only on morphisms between elements from $A^{1}$.
-Now let us recall the definition of an isomorphism with respect to the structure of DG–algebra.
+    Let $g\in A^{i}$, then $d(g)=\sum\limits_{h\in G}\chi(h,g)h\in A^{i+1}\Leftrightarrow\chi(h,g)=0$ when $h\notin A^{i+1}$.
+    Note that if $u\in A^{i+1},v\in A^{i}$, then for $\phi=(u,v)$ we have: $s(\phi)=(-1)^{|v|}v^{-1}u\in A^{1}$, $t(\phi)=uv^{-1}\in A^{1}$,which means that in the case of a DG–algebra, the characters may not be equal to 0 only on morphisms between elements from $A^{1}$.
+Now let us recall the definition of an isomorphism of DG–algebras.
 Definition 3.2.
-An isomorphism $f$ between two DG–algebras $(A_{1},d_{1})$ and $(A_{2},d_{2})$, $f$: $A_{1}$ $\to$ $A_{2}$, is an algebra isomorphism such that following properties hold:
+An isomorphism $f$: $A_{1}$ $\to$ $A_{2}$ between two DG–algebras $(A_{1},d_{1})$ and $(A_{2},d_{2})$ is an algebra isomorphism such that following properties hold:
   1. 1.
-    f holds the structure of graded algebra $f(A^{i}_{1})\subset A^{i}_{2}$.
+    f respects the grading, $f(A^{i}_{1})\subset A^{i}_{2}$;
   2. 2.
-    for all $a\in A_{1}$ holds $f(d_{1}(a))=d_{2}(f(a))$.
-Now we will describe the structures of DG–algebras up to isomorphism from definition (3.2).
-Let $X_{i},i=1,2$ be a matrix (if $G$ is an infinite group, that this matrix is also infinite) in following sense
-$X:G\times G\to\mathbb{C},$
-$(x,h)\mapsto\chi(x,h).$
+    for all $a\in A_{1}$, $f(d_{1}(a))=d_{2}(f(a))$.
+Now we will describe the structures of DG–algebras up to isomorphism per definition (3.2).
+Let $X_{i},i=1,2$ be two matrices (infinity ones, if $G$ is an infinite group) in following sense:
+$X_{i}:G\times G\to\mathbb{C},$
+$(x,h)\mapsto\chi_{i}(x,h).$
 Theorem 3.2.
-Two DG–algebras $(\mathbb{C}[G],d_{1})$ and $(\mathbb{C}[G],d_{2})$ are isomorphic iff there is a matrix C, such that
+Two DG–algebras $(\mathbb{C}[G],d_{1})$ and $(\mathbb{C}[G],d_{2})$ are isomorphic iff there is a matrix $C$ such that
   1. 1.
-    matrixes $X_{1}$ and $X_{2}$ are conjugated
-    $CX_{2}=X_{1}C,$
+    matrices $X_{1}$ and $X_{2}$ are conjugate, $CX_{2}=X_{1}C;$
   2. 2.
-    The mapping f is $Aut(\mathbb{C}[G])$.
-In other words that means that sets of constants $c_{h}^{g}$, gives $f(g)=\sum\limits_{h\in G}c_{h}^{g}h\in Aut(\mathbb{C}[G])$ and
+    the mapping f is in $\operatorname{Aut}(\mathbb{C}[G])$.
+In other words, that means that sets of constants $c_{h}^{g}$, gives $f(g)=\sum\limits_{h\in G}c_{h}^{g}h\in\operatorname{Aut}(\mathbb{C}[G])$ and
 $\sum\limits_{h\in G}c_{h}^{g}\chi_{2}(x,h)=\sum\limits_{h\in G}\chi_{1}(h,g)c_{x}^{h},\forall g,x\in G,$ (8)
@@
 Proof.
-Suppose that we have $f:(\mathbb{C}[G],d_{1})\rightarrow(\mathbb{C}[G],d_{2})$ — an isomorphism of DG–algebras given on generators by the formula $f(g)=\sum\limits_{h\in G}c_{h}^{g}h$.
-By assumption we have:
+Suppose that we have an isomorphism of DG–algebras $f:(\mathbb{C}[G],d_{1})\rightarrow(\mathbb{C}[G],d_{2})$ ,given on generators by the formula $f(g)=\sum\limits_{h\in G}c_{h}^{g}h$.
+By definition, we have
 $f(d_{1}(g))=d_{2}(f(g))$.
 Then
-$f(d_{1}(g))=f(\sum\limits_{h\in G}\chi_{1}(h,g)h)=\sum\limits_{h\in G}\chi_{1}(h,g)f(h)=\sum\limits_{h\in G}\chi_{1}(h,g)(\sum\limits_{x\in G}c_{x}^{h}x)=(\sum\limits_{h\in G}\chi_{1}(h,g)c_{x}^{h})x$;
-$d_{2}(f(g))=d_{2}(\sum\limits_{h\in G}c_{h}^{g}h)=\sum\limits_{h\in G}c_{h}^{g}d_{2}(h)=\sum\limits_{h\in G}c_{h}^{g}(\sum\limits_{x\in G}\chi_{2}(x,h)x)=(\sum\limits_{h\in G}c_{h}^{g}\chi_{2}(x,h))x$.
-From where we get our equality (8). To prove the converse, just define $f(g)=\sum\limits_{h\in G}c_{h}^{g}h$ and for the same reasons, we get the necessary equalities. ∎
+$f(d_{1}(g))=f\Big{(}\sum\limits_{h\in G}\chi_{1}(h,g)h\Big{)}=\sum\limits_{h\in G}\chi_{1}(h,g)f(h)=\sum\limits_{h\in G}\chi_{1}(h,g)\big{(}\sum\limits_{x\in G}c_{x}^{h}x\Big{)}=\Big{(}\sum\limits_{h\in G}\chi_{1}(h,g)c_{x}^{h}\Big{)}x$;
+$d_{2}(f(g))=d_{2}\Big{(}\sum\limits_{h\in G}c_{h}^{g}h\Big{)}=\sum\limits_{h\in G}c_{h}^{g}d_{2}(h)=\sum\limits_{h\in G}c_{h}^{g}(\sum\limits_{x\in G}\chi_{2}(x,h)x)=\Big{(}\sum\limits_{h\in G}c_{h}^{g}\chi_{2}(x,h)\Big{)}x$.
+Whence we get our equality (8). To prove the converse, just define $f(g)=\sum\limits_{h\in G}c_{h}^{g}h$ and for the same reasons, get the necessary equalities. ∎
 Corollary 3.1.
-In the case when $f\in Aut(G)$ equality (8) takes the form
+In the case when $f\in\operatorname{Aut(G)}$ the equality (8) takes the form
 $\chi_{1}(h,g)=\chi_{2}(f(h),f(g)).$
@@
 Example 3.1.
-In group algebra $\mathbb{C}[G]\cong\bigoplus A^{i}$, such that $A^{1}\neq 0$ and $\exists a\in Z(G)$ and $a\in A^{1}$, the inner derivation generated by the element $a$ according to the formula from proposition 4.1 defines the structure of the DG–algebra on $\mathbb{C}[G]$.
-Note that identity element of group lies in $A^{0}$.
+In a group algebra $\mathbb{C}[G]\cong\bigoplus A^{i}$ such that $A^{1}\neq 0$, and given $a\in Z(G)\cap A^{1}$, the inner derivation generated by the element $a$ according to the formula from proposition 4.1 defines a DG–algebra structure on $\mathbb{C}[G]$.
+Note that the identity element of the group lies in $A^{0}$.
 Proof.
-It must be shown that $d^{2}=0$ is fulfilled for such derivations. We know from Theorem (3.1) that this condition is equivalent to
-$\sum\limits_{h\in G}\chi(h,g)\chi(x,h)=0$, $\forall g,x\in G.$
-Consider the product $\chi(h,g)\chi(x,h)$, by definition of inner derivations, this product is not equal to 0 if and only if $(h,g)$ and $(x,h)$ $\in Hom(a,-a)$ or Hom(a,a)Hom(-a,a).
-Let $s(h,g)=s(x,h)=a$ and t(h,g)=t(x,h)=at(h,g)=t(x,h)=-a, then $(-1)^{|g|}g^{-1}h=(-1)^{|h|}h^{-1}x$ and $hg^{-1}=xh^{-1}$, but it is impossible because $a\in A^{1}\Rightarrow|h|\neq|g|.$
-Let $s(h,g)=t(x,h)=a$ and t(h,g)=s(x,h)=at(h,g)=s(x,h)=-a, then $(-1)^{|g|}g^{-1}h=xh^{-1}$ and $hg^{-1}=(-1)^{|h|}h^{-1}x,$ but this is similarly impossible because $|h|\neq|g|.$
+We have to show that $d^{2}=0$ for such a derivation. We know from Theorem (3.1) that this condition is equivalent to
+$\sum\limits_{h\in G}\chi(h,g)\chi(x,h)=0$; $\forall g,x\in G.$
+Consider the product $\chi(h,g)\chi(x,h)$. By the definition of an inner derivation, this product is not equal to 0 if and only if $(h,g)$ and $(x,h)$ $\in Hom(a,-a)$ or Hom(a,a)Hom(-a,a).
+Let $s(h,g)=s(x,h)=a$ and t(h,g)=t(x,h)=at(h,g)=t(x,h)=-a, then $(-1)^{|g|}g^{-1}h=(-1)^{|h|}h^{-1}x$ and $hg^{-1}=xh^{-1}$; but this is impossible because $a\in A^{1}\Rightarrow|h|\neq|g|.$
+Let $s(h,g)=t(x,h)=a$ and t(h,g)=s(x,h)=at(h,g)=s(x,h)=-a, then $(-1)^{|g|}g^{-1}h=xh^{-1}$ and $hg^{-1}=(-1)^{|h|}h^{-1}x,$; but this is similarly impossible because $|h|\neq|g|.$
 The remaining two cases are similar. ∎
 Proposition 3.1.
-Let $d^{\tau}_{z}(g)=\tau(g)gz$ - the central derivation given by the element $z\in A^{1}$. It defines the structure of a DG–algebra if and only if $\tau(g)\tau(gz)=0,\forall g\in G$.
+Let $d^{\tau}_{z}(g)=\tau(g)gz$ be the central derivation given by the element $z\in A^{1}$. It defines a DG–algebra structure if and only if $\tau(g)\tau(gz)=0\;\forall g\in G$.
 Proof.
@@
 […]
-Sets the structure of the DG–algebra on $\mathbb{C}[G]$.
+defines a DG–algebra structure on $\mathbb{C}[G]$.
 […]
 Proof.
-To prove, it is necessary to check only the fulfillment of the graduated Leibniz rule.
-And to do this, we need to make sure that the given $\tau$ is really a group character and take advantage of Proposition (3.1).
-It is not difficult to do this by simply going through all possible options.
+It is only necessary to check that the graded Leibniz rule is satisfied.
+To do this, we can make sure that the given $\tau$ is indeed a group character and then take advantage of Proposition (3.1).
+It is not difficult to do this by simply going through all the possibilities.
 ∎
 Corollary 3.1.
-In group graded algebra $\mathbb{C}[Z]\cong\bigoplus\mathbb{C}[x^{k}]$, where $\mathbb{Z}=\left\langle{x}\right\rangle$, it is possible to define the structure of the DG algebra using central derivation $d^{\tau}_{x}$, where
+In the graded group algebra $\mathbb{C}[\mathbb{Z}]\cong\bigoplus\mathbb{C}[x^{k}]$, where $\mathbb{Z}=\left\langle{x}\right\rangle$, it is possible to define a DG–algebra structure using the central derivation $d^{\tau}_{x}$, where
 […]
 Corollary 3.2.
-Consider the Heisenberg integer group of 3×3 upper triangular matrices
+Consider the Heisenberg integer group of 3×3 upper–triangular matrices
 $H_{3}=\left\langle{x,y,z\mid z=xyx^{-1}y^{-1},xz=zx,yz=zy}\right\rangle$.
 It is well known that any element of this group can be represented as $y^{b}z^{c}x^{a}$.
-Then $f(y^{b}z^{c}x^{a})=c\in\mathbb{Z}$ - an epimorphism into a group of integers.
-From here we get the graduation $\mathbb{C}[H_{3}]\cong\bigoplus_{k\in\mathbb{Z}}\mathbb{C}[z^{k}\left\langle{x,y}\right\rangle]$.
-And the central differentiation $d^{\tau}_{z}$, constructed similarly to Example 5.2, gives the structure of the DG–algebra on $\mathbb{C}[H_{3}]$.
+Then $f(y^{b}z^{c}x^{a})=c\in\mathbb{Z}$ is an epimorphism into a group of integers.
+From here we get the grading $\mathbb{C}[H_{3}]\cong\bigoplus_{k\in\mathbb{Z}}\mathbb{C}[z^{k}\left\langle{x,y}\right\rangle]$.
+And the central derivation $d^{\tau}_{z}$, constructed similarly to Example 5.2, gives a DG–algebra structure on $\mathbb{C}[H_{3}]$.